OpenStudy (anonymous):

find the first ten terms of the series 1-3/2+9/4

Parth (parthkohli):

\(\Large \color{midnightblue}{\rightarrow ({- 3 \over 2})^{n} }\) This is the formula :)

OpenStudy (campbell_st):

this is an alternating geometric sequence the ratio of the terms will give the common ratio use \[\frac{T _{2}}{T _{1}}=\frac{T _{3}}{T _{2}} = ...\] find the common ratio and then its previous term multiplied by the common ratio.... \[(\frac{-3}{2})^n\] works for n = 0, 1, 2, 3, ....... or it could be written \[(\frac{-3}{2})^{n-1}\] where n = 1, 2, 3, 4, 5,....